https://sites.google.com/a/pessto.org/wiki/pessto-wiki/home/night-reports/night-report-2018-02-04

Night Report 2018-02-04

Observers: Regis Cartier, Alessandro Razza 
TAT: Janet Chen, Michel Dennefeld
TO: Javier Alarcon (day), Mauricio Martinez (night)

Support team: QUB

Calibrations:  SOFI specflat, SpecArc, and FlatJHK.

Weather : Telescope closed at twilight because of thick clouds, we opened a few minutes after twilight. 40-60% of the sky covered. 

UT            Object
---------------------------------------------------------------------------------------------------------------------------------------------------------------
00:43:00  Hip28112 non-centred on the central pixel.
01:02:36  Hip39864 aborted. Not visible on the acquisition, most likely behind thick clouds.
01:03:50  Hip21024 not-centred on the central pixel.
01:27:08  Hip21024 repeated now centred.
01:32:52  SN2017ilo JHK imaging seeing ~1.2.
02:04:30  SN2017jfa JH imaging. seeing~1.2.
02:52:51  OGLEjei JHK imaging seeing 1.0-0.7, airmass 1.6-1.7.

03:21:00  Switch to EFOSC to do classifications.

03:49:18  AT2018jm seeing~0.8  SN Ia; z=0.065 at +12 d. Classification.
04:00:26  AT2018lm seeing~0.9 SN Ia;  z=0.044 at max. Classification.
04:11:42  STD_LTT3864 seeing~1.0.
04:31:40  AT2018ly seeing ~ 1.1 CV? Classification.

04:37:30  Switch to SOFI to complete the NIR observing requests.

05:00:25  STD_LTT3218 flux standard observed in BG. Some problems with SOFI electronics. 
05:19:19  Problems with SOFI  while centring SN 2018ec on the slit, we lost 20 min.
05:40:54  SN2018ec BG spectrum seeing~0.8.
07:33:02  Telluric star HIP50489 in BG seeing~1.44
07:3818   SN2018ec JHKs imaging (less 90 deg)
08:09:09  SN2017jfv JHKs imaging on source after a few images aborted because of clouds, we will try tomorrow again.
08:27:16  STD_LTT3218 flux standard observed in BG. No problems with SOFI electronics, but observed through clouds.
08:39:34  Hip39864 observed through clouds.
08:46:09  Hip58316 observed through clouds.
08:53:44  Hip58316 observed through clouds and we change the exposure time to test the OB

Completely covered at the end, but it was a productive night.






Comments